# From Kernels to Attention: A Transformer Framework for Density and Score Estimation

### v1

@@ Abstract
We introduce a unified attention-based framework for joint score and density estimation. Framing the problem as a sequence-to-sequence task, we develop a permutation- and affine-equivariant transformer that estimates both the probability density $f(x)$ and its score $\nabla_{x}\log f(x)$ directly from i.i.d. samples. Unlike traditional score-matching methods that require training a separate model for each distribution, our approach learns a single distribution-agnostic operator that generalizes across densities and sample sizes. The architecture employs cross-attention to connect observed samples with arbitrary query points, enabling generalization beyond the training data, while built-in symmetry constraints ensure equivariance to permutation and affine transformations. Analytically, we show that the attention weights can recover classical kernel density estimation (KDE), and verify it empirically, establishing a principled link between classical KDE and the transformer architecture. Empirically, the model achieves substantially lower error and better scaling than KDE and score-debiased KDE (SD-KDE), while exhibiting better runtime scaling. Together, these results establish transformers as general-purpose, data-adaptive operators for nonparametric density and score estimation.
Machine Learning, ICML
Vasily Ilin1Peter Sushko2
1Department of Mathematics, University of Washington, Seattle, USA
2Allen Institute for Artificial Intelligence, Seattle, USA
@@ 1 Introduction
Estimating a probability density $f$ and its score $\nabla\log f$ from i.i.d. samples underpins inference, simulation, and control in modern machine learning and statistics. Classical nonparametric estimators such as kernel density estimation (KDE) face a sharp bias–variance trade-off and deteriorate in moderate dimensions. Score-based methods alleviate some issues but typically learn a model per target density, limiting reuse across datasets and tasks. While many classical estimators (including KDE) respect permutation and affine transformations at the level of their functional definitions, modern learned approaches rarely build in these symmetries architecturally, leaving density and score estimates vulnerable to unnecessary sample- and feature-space dependence.
We consider the task of estimating both $f(x)$ and $\nabla_{x}\log f(x)$ from an i.i.d. sample $X\in\mathbb{R}^{n\times d}$ of $f$:
$\begin{pmatrix}x_{1}\\[-2.0pt] \vdots\\[-2.0pt] x_{n}\end{pmatrix}\xrightarrow{\;T\;}\begin{pmatrix}f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] f(x_{n})\end{pmatrix},\qquad\begin{pmatrix}x_{1}\\[-2.0pt] \vdots\\[-2.0pt] x_{n}\end{pmatrix}\xrightarrow{\;S\;}\begin{pmatrix}\nabla\log f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla\log f(x_{n})\end{pmatrix}$.
We recognize this as a sequence-to-sequence modeling task (Sutskever et al., 2014) subject to two key symmetries acting on the index and feature spaces:
$T(PXA+\mu)$ $=P\,|\det A|^{-1}\,T(X),$ (1)
$S(PXA+\mu)$ $=P\,(A^{\top})^{-1}\,S(X),$ (2)
for a permutation matrix $P\in\mathbb{R}^{n\times n}$, invertible matrix $A\in\mathbb{R}^{d\times d}$, and shift $\mu\in\mathbb{R}^{d}$. These relations express how density and score estimates must transform under permutation and affine transformations of the sample.
Building on this observation, we propose a permutation- and affine-equivariant Transformer that learns these mappings directly from data. The model treats the sample $X$ as a sequence and is thus permutation-equivariant by construction. To achieve affine equivariance, we introduce a lightweight whitening–rewrapping mechanism in the forward pass, ensuring the outputs transform exactly as densities and scores should under affine changes of coordinates. A single network with two output heads jointly predicts $f(x)$ and $\nabla\log f(x)$, supporting variable sample sizes $n$ and feature dimensions $d$.
Analytically, we show that under suitable parameterization, the self-attention weights recover normalized KDE weights, establishing a direct link between classical kernel averaging and learned attention. This connection provides a principled interpretation of the Transformer’s smoothing behavior and grounds its ability to adapt bandwidths and kernels from data.
Trained on synthetic Gaussian mixtures generated on the fly, the model learns distribution-agnostic rules that generalize to unseen dimensions, sample sizes, and mixture complexities. Empirically, it outperforms KDE and score-debiased KDE across both density and score estimation tasks, while exhibiting favorable runtime and memory scaling. Beyond estimation, the learned density and score serve as accurate plug-in oracles for downstream computations such as entropy, Fisher information, and as components in solvers for Fokker–Planck–type PDEs.
Together, these results demonstrate that symmetry-respecting Transformers can serve as general-purpose operators for nonparametric density and score estimation, combining the interpretability of classical estimators with the flexibility and scalability of modern deep learning.
@@
  • A permutation- and affine-equivariant Transformer that jointly predicts density and score from i.i.d. samples and generalizes across $n$ and $d$.
  • A theoretical link showing self-attention can implement normalized KDE weights, yielding a principled, data-adaptive smoother.
  • A cross-attention formulation for evaluating $f$ and $\nabla\log f$ at arbitrary query sets without retraining.
  • Extensive experiments demonstrating improved accuracy and scaling vs. KDE and score-debiased KDE, plus effective plug-in estimates for entropy/Fisher information and strong performance in kinetic PDE solvers.
@@ 2 Related Work
Our setting lies at the intersection of operator learning for probability measures, transformer-based modeling of permutation-invariant data, and score-assisted nonparametric density estimation.
@@
Classical nonparametric density estimation methods such as Parzen windows (Parzen, 1962) and Silverman’s rule-of-thumb KDE (Silverman, 2018; Wand & Jones, 1994) rely on local averaging with a fixed bandwidth, suffering from the usual bias–variance trade-off and poor scaling in high dimensions. Score-Debiased KDE (SD-KDE) (Epstein et al., 2025) shows that if a score oracle is available, one can remove the leading-order bias of KDE by a single correction step. Our model provides precisely this missing ingredient: a reusable, density-agnostic score oracle that can plug into SD-KDE or related estimators.
Transformers for permutation-invariant and mean-field systems.
Permutation-equivariant architectures such as DeepSets (Zaheer et al., 2017), Set Transformers (Lee et al., 2019), and equivariant graph networks (Maron et al., 2018) capture exchangeable structure in sets and particles. More recently, permutation-equivariant transformers have been shown to approximate vector fields of interacting-particle and mean-field systems (Biswal et al., 2024a, b). Our model exploits the same symmetry but outputs statistical quantities—density and score—while additionally enforcing affine equivariance.
@@
Neural Operator and Fourier Operator frameworks (Kovachki et al., 2023) learn mappings between infinite-dimensional function spaces, and the Score Neural Operator (Liao et al., 2024) extends this idea to distributions via RKHS embeddings. In contrast, our transformer acts directly on raw i.i.d. samples, avoiding RKHS representations and yielding a simple symmetry-aware operator that generalizes across densities and dimensions.
Transformers for Gaussian mixtures.
TGMM (Chen et al., 2025) and related models learn to infer or reconstruct parameters of Gaussian mixtures. Such parametric approaches are inherently limited to mixture families, whereas our method is fully nonparametric and distribution-agnostic, producing pointwise density and score estimates for arbitrary smooth targets.
Score-based and diffusion models.
Classical score matching (Hyvärinen, 2005) and its denoising and diffusion extensions (Song & Ermon, 2019; Song et al., 2021a, b) train models to recover the score of a single density along a diffusion process. Our approach instead learns one universal, permutation- and affine-equivariant network that generalizes across densities without retraining, providing a fast drop-in oracle for downstream estimators such as Fisher information, entropy, and Fokker–Planck solvers.
@@ 3 Methodology
3.1 Inductive Bias
@@
Let $P\in\mathbb{R}^{n\times n}$ be a permutation matrix, $A\in\mathbb{R}^{d\times d}$ be invertible, and $\mu\in\mathbb{R}^{d}$. Then
$T\big(PXA+\mu\big)$ $=P\,|\det A|^{-1}\,T(X),$ (3)
$S\big(PXA+\mu\big)$ $=P\,(A^{T})^{-1}\,S(X).$ (4)
To capture the permutation and affine equivariance, we use the Transformer architecture without positional encodings, and with affine normalization, as shown in Figure 1.
def forward(self, X, Y):
@@
return dens / torch.det(A), score @ A.T
@@
Additionally, Proposition 3.2 demonstrates that a single attention head can compute the classical KDE. This shows that the Transformer architecture is well-suited to the density estimation task.
Proposition 3.2 (Attention computes KDE).
@@
While the above Proposition establishes that the Attention mechanism has the capacity to approximate KDE well, we must verify whether the model learns to do it practice. To empirically verify the relationship between the KDE weights and attention scores, we visualize attention in Figure 2. As expected, the Transformer learns to attend to nearby points and disregard far away samples. The attention mechanism effectively learns a kernel-like behavior, with emergent head specialization. The learned features allow for much better scaling both in the sample size and dimension, compared to KDE. In section B of the Appendix we further visualize the attention scores of individual heads, and find an emergent behavior of head specialization. Specifically, head 1 specialized to look at far-away points, whereas heads 0, 2, and 5 specialized to look at close- and mid-range interactions. Finally, heads 3, 4, 6, and 7 specialized to look in specific directions. This emergent behavior further links the multi-head Transformer to kernel-based methods, but with flexible multi-scale kernels.
@@ 3.2 Training Data
Input: batch size $B$, dimension $d$, sample sizes $n_{x},n_{y}$, components $k_{\min},k_{\max}]$, mean scale $\mu$, cov noise $\varepsilon$
@@
   Sample weights: $w_{b}\in\mathbb{R}^{k}_{+}$
   Sample means: $\mu_{b,j}\sim\mathrm{Uniform}([-\mu,\mu]^{d})$
   Sample covs: $A\sim\mathcal{N}(0,I_{d})$ and $\Sigma_{b,j}=AA^{\top}+\varepsilon I_{d}$
   Sample $X_{b}\sim\mathrm{GMM}_{b}$, and $Y_{b}\sim\mathrm{GMM}_{b}$
   Densities: $d_{b}(y)=\sum_{j}w_{b,j}\,\mathcal{N}(y\mid\mu_{b,j},\Sigma_{b,j})$
   Scores: $s_{b}(y)=\nabla_{y}\log d_{b}(y)$ for all $y\in Y_{b}$
@@
3.3 Models Variants and Implementation
@@
Recognizing that density and score estimation are mathematically coupled—linked through the gradient of $\log f$—we employ a unified architecture with a shared backbone and two specialized output heads. This joint formulation encourages mutual consistency between the two quantities and allows the model to leverage shared geometric and statistical features during training, improving both accuracy and sample efficiency.
To further enhance generality, we train a single model capable of handling inputs of varying dimensionality, from one to five dimensions. Samples are embedded into a fixed-size tensor representation through a lightweight zero-padding scheme, preserving numerical stability while maintaining permutation and affine equivariance. This design choice enables seamless transfer of learned representations across different dimensional settings and supports robust estimation in diverse regimes of $n$ and $d$.
@@ 4 Experiments
In this section we empirically answer the following questions:
@@
    How does our model compare to KDE and its score-debiased variant SD-KDE (Epstein et al., 2025)?
@@
    How does our model scale in the sample size $n$ and dimension $d$?
@@
    How well does our model generalize to out-of-distribution data, such as GMMs with more modes?
@@
    How does the inference runtime scale in $n$ and $d$?
We also show that the learned model can serve as a score oracle in SD-KDE, as well as deterministic SBTM-like solvers for Fokker-Planck-type equations such as the Landau equation.
@@
All experiments are performed on a single 48GB L40S GPU. The architecture is a permutation- and affine-equivariant Transformer. It consists of a linear embedding, four Transformer encoder layers (128-dimensional hidden size, 8 heads, GELU activation, pre-normalization), and a linear projection to the output dimension. No positional encodings are used, preserving permutation-equivariance. The model has about $800,000$ parameters. We use batch size $32$, sample size $n=2048$, dropout $0.1$, and draw GMMs with 1-10 modes with means in $3,3]^{d}$ and diagonal covariances in $0.2,1]^{d}$. We use learning rate $0.0001$ for low dimensions and $0.001$ for higher dimensions.

### v2

# DiScoFormer: Plug-In Density and Score Estimation with Transformers

@@ Abstract
Estimating probability density and its score from samples remains a core problem in generative modeling, Bayesian inference, and kinetic theory. Existing methods are bifurcated: classical kernel density estimators (KDE) generalize across distributions but suffer from the curse of dimensionality, while modern neural score models achieve high precision but require retraining for every target distribution. We introduce DiScoFormer (Density and Score Transformer), a “train-once, infer-anywhere” equivariant Transformer that maps i.i.d. samples to both density values and score vectors, generalizing across distributions and sample sizes. Analytically, we prove that self-attention can recover normalized KDE, establishing it as a functional generalization of kernel methods; empirically, individual attention heads learn multi-scale, kernel-like behaviors. The model converges faster and achieves higher precision than KDE for density estimation, and provides a high-fidelity plug-in score oracle for score-debiased KDE, Fisher information computation, and Fokker-Planck-type PDEs.
density estimation,score estimation,transformers,kernel methods,equivariant networks,nonparametric statistics
@@ 1 Introduction
Estimating a probability density $f$ and its corresponding score function $\nabla\log f$ from i.i.d. samples is a foundational problem in statistical inference, underpinning tasks in generative modeling, stochastic control, and information theory. While classical non-parametric methods, such as Kernel Density Estimation (KDE), offer strong theoretical guarantees and respect fundamental symmetries, they are hampered by a rigid bias-variance trade-off and a prohibitive “curse of dimensionality” as dimension $d$ increases. Conversely, modern neural score-matching models achieve high accuracy in high-dimensional settings but are typically transductive: they must be retrained for every new target distribution, precluding their use as general-purpose, “off-the-shelf” estimators.
We propose bridging this gap by framing density and score estimation as a sequence-to-operator (Sutskever et al., 2014) learning task. Given an i.i.d. sample $X=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$, we seek to learn a single model that maps the entire sample $X$ to its underlying density and score functions. Specifically, we define an operator $T$ for density and $S$ for scores:
$\begin{pmatrix}x_{1}\\[-2.0pt] \vdots\\[-2.0pt] x_{n}\end{pmatrix}\xrightarrow{\;T\;}\begin{pmatrix}f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] f(x_{n})\end{pmatrix},\qquad\begin{pmatrix}x_{1}\\[-2.0pt] \vdots\\[-2.0pt] x_{n}\end{pmatrix}\xrightarrow{\;S\;}\begin{pmatrix}\nabla\log f(x_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla\log f(x_{n})\end{pmatrix}$. (1)
A critical requirement for such an operator is that it must respect the inherent symmetries of the data. Specifically, the estimates should be permutation-equivariant with respect to the sample index and affine-equivariant with respect to the coordinate space. For a permutation matrix $P$, invertible matrix $A$, and shift $\mu$, these symmetries are defined as:
$T(PXA+\mathbf{1}\mu^{\top})$ $=P\,|\det A|^{-1}\,T(X),$ (2)
$S(PXA+\mathbf{1}\mu^{\top})$ $=P\,A^{-1}\,S(X),$ (3)
Building on these principles, we introduce DiScoFormer – Density and Score Transformer. By treating the sample $X$ as a sequence, the Transformer architecture provides permutation equivariance by construction. We achieve affine equivariance through a specialized whitening mechanism that normalizes inputs up to orthogonal transformations, combined with data augmentation to ensure rotation invariance. Unlike traditional models, our architecture utilizes cross-attention, allowing for the evaluation of $f$ and $\nabla\log f$ at arbitrary query points without retraining.
Furthermore, we provide a theoretical bridge between modern attention mechanisms and classical non-parametrics. We demonstrate analytically that under specific parameterizations, self-attention weights recover normalized KDE weights. This expressivity result positions the Transformer not as a black box, but as a principled, data-adaptive generalization of kernel methods that can dynamically learn multiscale smoothing behaviors from data.
Empirically, the model outperforms KDE and score-debiased KDE (Epstein et al., 2025) across a range of dimensions and sample sizes on both in-distribution and out-of-distribution test cases. Trained on Gaussian Mixture Model (GMM) data, the model generalizes to GMMs with more modes and non-Gaussian targets such as the Laplace distribution and demonstrates favorable scaling in both $n$ and $d$. By amortizing the estimation cost, our model provides high-fidelity plug-in oracles for downstream tasks, including entropy estimation, Fisher information calculation, and solving Fokker-Planck-type partial differential equations (PDEs).
@@
  • DiScoFormer, a universal Transformer model for one-shot estimation of density and score functions from i.i.d. samples.
  • A proof and empirical evidence establishing self-attention as a data-adaptive generalization of normalized kernel density estimation.
  • Non-parametric density and score estimators that surpass classical KDE methods in accuracy across sample sizes and dimensions.
  • Applications to Fisher information, differential entropy, and Fokker-Planck-type PDEs.
@@ 2 Related Work
Our work lies at the intersection of nonparametric density estimation, score-based methods, permutation-invariant neural architectures, and operator learning for probability measures.
@@
Classical nonparametric methods such as Parzen windows (Parzen, 1962) and Silverman’s rule-of-thumb KDE (Silverman, 1986; Wand and Jones, 1994) estimate densities via local kernel averaging. While theoretically grounded, these methods suffer from a rigid bias–variance trade-off and scale poorly with dimension (Scott, 2015). Adaptive bandwidth methods (Abramson, 1982) and fast algorithms such as the Fast Gauss Transform (Greengard and Strain, 1991) partially alleviate computational costs but do not resolve the fundamental curse of dimensionality. Score-Debiased KDE (Epstein et al., 2025) demonstrates that access to a score oracle enables bias correction, motivating our approach: we provide precisely this missing oracle as a reusable, distribution-agnostic component.
Other nonparametric density estimators.
$k$-nearest-neighbor methods (Loftsgaarden and Quesenberry, 1965), orthogonal series estimators (Efromovich, 2010; Scott, 1985), and normalizing flows (Rezende and Mohamed, 2015; Dinh et al., 2017; Papamakarios et al., 2017) offer alternatives to KDE but either lack direct score estimates or require per-distribution retraining. Our model provides both density and score from a single pretrained network.
Score matching and score-based generative models.
Score matching (Hyvärinen, 2005) and its denoising variant (Vincent, 2011) estimate unnormalized model scores without computing partition functions. Sliced score matching (Song et al., 2019) scales these ideas to higher dimensions. Score-based generative models (Song and Ermon, 2019; Ho et al., 2020; Song et al., 2021b, a) learn scores along diffusion processes for sampling, with fast solvers (Lu et al., 2022; Karras et al., 2022) and convergence guarantees (Chen et al., 2023, 2024). These methods train on a single target density; in contrast, our model learns a universal score operator that generalizes across distributions without retraining.
Permutation-invariant architectures.
DeepSets (Zaheer et al., 2017) and Set Transformers (Lee et al., 2019) provide foundational architectures for exchangeable data, while equivariant graph networks (Maron et al., 2019) extend these ideas to richer symmetry groups. Neural Processes (Garnelo et al., 2018; Edwards and Storkey, 2017) and Perceiver IO (Jaegle et al., 2022) learn distributions over functions from context sets. Recent work shows that permutation-equivariant transformers can universally approximate mean-field dynamics (Biswal et al., 2025). Our architecture exploits the same symmetry but outputs statistical quantities—density and score—while additionally enforcing affine equivariance.
Attention as kernel smoothing.
Formal connections between softmax attention and kernel methods have been developed through linear-attention reformulations (Katharopoulos et al., 2020), random-feature approximations (Choromanski et al., 2021), Nyström methods (Xiong et al., 2021), explicit Gaussian-kernel replacements (Chen et al., 2021; Lu et al., 2021a), and Nadaraya–Watson regression interpretations (Zhang et al., 2023). Our Proposition 3.2 builds on this lineage by showing that self-attention weights can exactly reproduce normalized Gaussian KDE weights.
@@
Neural Operators (Kovachki et al., 2023), DeepONet (Lu et al., 2021b), and Fourier Neural Operators (Li et al., 2021) learn mappings between infinite-dimensional function spaces. The Score Neural Operator (Liao et al., 2024) extends this framework to probability distributions via RKHS embeddings. In contrast, our transformer acts directly on raw i.i.d. samples without kernel embeddings, yielding a simpler operator that generalizes across densities and dimensions.
Particle methods and Fokker–Planck solvers.
Interacting particle methods for Fokker–Planck equations (Carrillo et al., 2019; Maoutsa et al., 2020) and score-based transport modeling (Boffi and Vanden-Eijnden, 2023; Lu et al., 2024; Ilin et al., 2025a) require accurate score estimates at each timestep. Stein Variational Gradient Descent (Liu and Wang, 2016; Liu, 2017; Duncan et al., 2023) similarly relies on kernel-based score approximations. Our model provides a fast, pretrained oracle that plugs directly into these solvers, avoiding per-distribution retraining.
@@ 3 Methodology
3.1 Equivariance
Table 1: Relative MSE of affine equivariance error, averaged over 50 trials.
Transform Rel. MSE
Permutation 0
Translation 0
Isotropic scaling 0
Anisotropic scaling 0
Rotation $5\times 10^{-4}$
Full affine $1\times 10^{-4}$
@@ 3.1 Equivariance
Let $P\in\mathbb{R}^{n\times n}$ be a permutation matrix, $A\in\mathbb{R}^{d\times d}$ be invertible, $\mu\in\mathbb{R}^{d}$, and $\mathbf{1}\in\mathbb{R}^{n}$ be the vector of ones. Then
$T\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,|\det A|^{-1}\,T(X),$ (4)
$S\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,A^{-1}\,S(X).$ (5)
To capture these symmetries, we use the Transformer architecture without positional encodings, combined with an affine normalization layer (Figure 2). The whitening step yields equivariance up to rigid rotations $O(d)$; training on randomly oriented GMMs encourages the network to learn approximate rotation invariance, closing the remaining gap in practice. Table 1 confirms this empirically.
@@
return dens * torch.det(A), score @ A.T
@@
3.2 Attention and KDE
Proposition 3.2 demonstrates that a single attention head can compute the classical KDE, showing that the attention mechanism has the capacity to implement kernel-based smoothing. While the proposition assumes $L^{2}$-normalized inputs, in practice the learned $Q$ and $K$ projections adapt to the actual input statistics after whitening.
Proposition 3.2 (Attention can compute KDE).
@@ 3.2 Attention and KDE
While the above Proposition establishes that the Attention mechanism has the capacity to approximate KDE well, we must verify whether the model learns to do it practice. To empirically verify the relationship between the KDE weights and attention scores, we visualize attention in Figure 1. As expected, the Transformer learns to attend to nearby points, with one attention head attending over far away samples instead. The attention mechanism effectively learns a kernel-like behavior, with emergent head specialization. The learned features allow for much better scaling both in the sample size and dimension, compared to KDE. In section B of the Appendix we further visualize the attention scores of individual heads, and find an emergent behavior of head specialization. Specifically, head 1 specialized to look at far-away points, whereas heads 0, 2, and 5 specialized to look at close- and mid-range interactions. Finally, heads 3, 4, 6, and 7 specialized to look in specific directions. This emergent behavior further links the multi-head Transformer to kernel-based methods, but with flexible multi-scale kernels.
@@ 3.3 Training Data
Input: batch size $B$, dimension $d$, sample sizes $n_{x},n_{y}$, components $k_{\min},k_{\max}]$
@@
   Sample two random GMMs with $k$ components each
   Sample $X_{b}$ from the first, $Y_{b}$ from the second
   Compute $f_{X_{b}}(y)$ and $\nabla\log f_{X_{b}}(y)$ for $y\in Y_{b}$
@@
3.4 Model Variants and Implementation
@@
Recognizing that density and score estimation are mathematically coupled – linked through the gradient of $\log f$ – we employ a unified architecture with a shared backbone and two specialized output heads. This joint formulation encourages mutual consistency between the two quantities and allows the model to leverage shared geometric and statistical features during training, improving both accuracy and sample efficiency.
@@ 4 Experiments
Figure 3: Score estimation comparison between Silverman KDE (Silverman, 1986) and our transformer model. The transformer is more accurate, especially in the sparse regions. We plot the negated score for easier viewing.
In this section, we empirically investigate:
@@
    comparison with KDE and SD-KDE (Epstein et al., 2025),
@@
    scaling in sample size $n$ and dimension $d$,
@@
    out-of-distribution generalization,
@@
    inference-time scaling in $n$ and $d$.
We also show that the learned model can serve as a score oracle in SD-KDE, as well as in deterministic SBTM-like solvers for Fokker-Planck-type equations such as the Landau equation.
@@
All experiments are performed on a single 48GB L40S GPU. The architecture is a permutation- and affine-equivariant Transformer. It consists of a linear embedding, four Transformer encoder layers (128-dimensional hidden size, 8 heads, GELU activation, pre-normalization), and a linear projection to the output dimension. No positional encodings are used, preserving permutation-equivariance. The model has about $800,000$ parameters. We use batch size $32$, sample size $n=2048$, dropout $0.1$, and draw GMMs with 1–10 modes with means in $3,3]^{d}$ and diagonal covariances in $0,1]^{d}$. A runtime comparison with KDE is provided in Appendix D.